# The separating Noether number of small groups

### v2

@@ 6. Some groups $G=H\times\mathrm{C}_{2}$ with $\beta_{\mathrm{sep}}^{K}(G)=\beta_{\mathrm{sep}}^{K}(H)$
6.1. The group $\mathrm{Dic}_{8}\times\mathrm{C}_{2}$
In this section
$G=\mathrm{Dic}_{8}\times\mathrm{C}_{2}=\langle a,b\mid a^{4}=1,\ b^{2}=a^{2},\ ba=a^{3}b\rangle\times\langle c\mid c^{2}=1\rangle$
is the direct product of the quaternion group of order $8$ and the group of order $2$. Assume that $K$ has an element $\mathrm{i}$ of multiplicative order $4$. Then $G$ has two irreducible $2$-dimensional representations (up to isomorphism), namely
$\psi_{1}:a\mapsto\begin{bmatrix}\mathrm{i}&0\\ 0&-\mathrm{i}\\ \end{bmatrix},\quad b\mapsto\begin{bmatrix}0&-1\\ 1&0\\ \end{bmatrix},\quad c\mapsto\begin{bmatrix}1&0\\ 0&1\\ \end{bmatrix}$
(this is the only $2$ dimensional representation of $\mathrm{Dic}_{8}$ lifted to $G$ by the projection from $G$ to its direct factor $\mathrm{C}_{2}$), and
$\psi_{2}:a\mapsto\begin{bmatrix}\mathrm{i}&0\\ 0&-\mathrm{i}\\ \end{bmatrix},\quad b\mapsto\begin{bmatrix}0&-1\\ 1&0\\ \end{bmatrix},\quad c\mapsto\begin{bmatrix}-1&0\\ 0&-1\\ \end{bmatrix},$
the representation $\psi_{1}$ tensored with the non-trivial representation $c\mapsto-1$ of $\mathrm{C}_{2}$. The other irreducible representations of $G$ are $1$-dimensional, and can be labeled by the elements of the subgroup $\widehat{G}:=\{\pm 1\}\times\{\pm 1\}\times\{\pm 1\}$ of $K^{\times}\times K^{\times}\times K^{\times}$ as follows: identify $\chi=(\chi_{1},\chi_{2},\chi_{3})\in\widehat{G}$ with the representation $\chi:G\to K^{\times}$ given by
$\chi:a\mapsto\chi_{1},\quad b\mapsto\chi_{2},\quad c\mapsto\chi_{3}.$
Denote by $W_{j}$ the vector space $K^{2}$ endowed with the representation $\psi_{j}$ for $j=1,2$, and for $\chi\in\widehat{G}$ denote by $U_{\chi}$ the vector space $K$ endowed with the representation $\chi$. Set
$U:=\bigoplus_{\chi\in\widehat{G}}U_{\chi}.$
The coordinate functions on $W_{1}$ are denoted by $x_{1},x_{2}$, the coordinate functions on $W_{2}$ are $y_{1},y_{2}$, and the coordinate function on $U_{\chi}$ is $t_{\chi}$ for $\chi\in\widehat{G}$ (see Section 3.6).
Proposition 6.1.
For $i=1,2$ we have $\beta(G,W_{i}\oplus U)=6$.
Proof.
First we deal with the case $i=1$. It is well known (see for example the proof of Proposition 4.2) that
(15) $K[W_{1}]^{G}=K[(x_{1}x_{2})^{2},\ x_{1}^{4}+x_{2}^{4},\ x_{1}x_{2}(x_{1}^{4}-x_{2}^{4})].$
It is easy to verify that a Gröbner basis with respect to the lexicographic term order induced by $x_{1}>x_{2}$ of the Hilbert ideal $\mathcal{H}(G,W_{1})$ is $x_{1}^{2}x_{2}^{2}$, $x_{1}^{4}+x_{2}^{4}$, $x_{2}^{6}$, $x_{1}x_{2}^{5}$. Using this Gröbner basis one can show that setting
$s_{(1,-1,1)}:=x_{1}x_{2},\quad s_{(-1,1,1)}:=x_{1}^{2}+x_{2}^{2},\quad s_{(-1,-1,1)}:=x_{1}^{2}-x_{2}^{2},$
a $G$-module direct complement in $K[x_{1},x_{2}]$ of the Hilbert ideal $\mathcal{H}(G,W_{1})$ is
$K\oplus\mathrm{Span}_{K}\{x_{1},x_{2}\}\oplus Ks_{(1,-1,1)}\oplus Ks_{(-1,1,1)}\oplus Ks_{(-1,-1,1)}\oplus\mathrm{Span}_{K}\{x_{1}^{3},x_{2}^{3}\}$
$\oplus\mathrm{Span}_{K}\{x_{1}^{2}x_{2},x_{1}x_{2}^{2}\}\oplus Ks_{(1,-1,1)}s_{(-1,1,1)}\oplus Ks_{(1,-1,1)}s_{(-1,-1,1)}$
$\oplus Ks_{(-1,1,1)}s_{(-1,-1,1)}\oplus\mathrm{Span}_{K}\{x_{1}^{5},x_{2}^{5}\}.$
The direct summands above are minimal $G$-invariant subspaces. Moreover, $G$ acts on $Ks_{\chi}$ via the character $\chi$. It follows by Lemma 3.9 that $K[W_{1}\oplus U]^{G}$ is generated over $K[W_{1}]^{G}$ by products of elements from $\{s_{\chi},\quad t_{\chi^{\prime}}\mid\chi\in\{(1,-1,1),(-1,1,1),(-1,-1,1)\},\ \chi^{\prime}\in\widehat{G}\}$, with at most two factors of the form $s_{\chi}$, and no $G$-invariant proper subproduct. Since the Davenport constant of the group $\widehat{G}$ is $4$, it is sufficient to consider products with at most $4$ factors. All of them have degree at most $6$.
The case $i=2$ is essentially the same. Indeed, the map $x_{1}\mapsto y_{1}$, $x_{2}\mapsto y_{2}$ extends to a $K$-algebra isomorphism between $K[W_{1}]$ and $K[W_{2}]$, which is also a $\mathrm{Dic}_{8}$-module isomorphism. The subgroup $\langle c\rangle$ acts trivially on $K[W_{1}]$, and acts by multiplication by $-1$ of the variables $y_{1},y_{2}$ in $K[W_{2}]$. Therefore replacing $x_{1}$ by $y_{1}$ and $x_{2}$ by $y_{2}$ in the corresponding formulae above, we get the generators of $K[W_{2}]^{G}$, as well as a direct sum decomposition into minimal $G$-invariant subspaces of a $G$-module direct complement in $K[W_{2}]$ of the Hilbert ideal. So in the same way as in the above paragraph, we get the conclusion $\beta(W_{2}\oplus U)=6$. ∎
Proposition 6.2.
We have $\beta(G,W_{1}\oplus W_{2})=6$.
Proof.
The element $a^{2}\in G$ multiplies each of the indeterminates $x_{1},x_{2},y_{1},y_{2}$ by $-1$. Since $G$-invariants are $a^{2}$-invariant, they must involve only monomials of even degree. Thus $\beta(G,W_{1}\oplus W_{2})$ is even. On the other hand, $\beta^{K}(G)=7$ by [15, Proposition 3.2]. We conclude that $\beta(G,W_{1}\oplus W_{2})\leq 6$. Note also that $\beta(G,W_{1}\oplus W_{2})\geq\beta(G,W_{1})$, and we saw in the course of the proof of Proposition 6.1 that $\beta(G,W_{1})=6$ (the third generator in (15) is not symmetric in $x_{1}$ and $x_{2}$, hence can not be expressed by the first two). ∎
Proposition 6.3.
Let $v:=(w_{1},w_{2})\in W_{1}\oplus W_{2}$ with $w_{1}\neq 0$ and $w_{2}\neq 0$. Then for any $\chi\in\widehat{G}$ there exists a relative invariant $f$ of weight $\chi$ such that $\deg(f)\leq 4$ and $f(v)\neq 0$.
Proof.
$\begin{array}[]{c|c}\chi&\text{relative invariants}\\ \hline\cr\hline\cr(-1,1,1)&f_{1}:=x_{1}^{2}+x_{2}^{2},\quad f_{2}:=x_{1}x_{2}(x_{1}^{2}-x_{2}^{2})\\ \hline\cr(-1,1,-1)&f_{1}:=x_{1}y_{1}+x_{2}y_{2},\quad f_{2}:=x_{1}x_{2}(x_{1}y_{1}-x_{2}y_{2}),\quad f_{3}:=x_{2}^{3}y_{1}-x_{1}^{3}y_{2}\\ \hline\cr(1,1,-1)&f_{1}:=x_{1}y_{2}-x_{2}y_{1},\quad f_{2}:=x_{1}x_{2}(x_{1}y_{2}+x_{2}y_{1}),\quad f_{3}:=x_{2}^{3}y_{2}+x_{1}^{3}y_{1}\end{array}$
A row in the above table contains a character $\chi\in\widehat{G}$ and relative invariants of weight $\chi$ of degree at most $4$, such that the common zero-set of these relative invariants is contained in the union of $W_{1}\oplus\{0\}$ and $W_{2}\oplus\{0\}$. For the other non-trivial weights the result can be deduced without further computation using Lemma 3.11. Indeed, the automorphism group of $G$ contains the automorphism group of its subgroup $\mathrm{Dic}_{8}$ as a subgroup. The automorphism group of $\mathrm{Dic}_{8}$ acts (on the right) therefore on $\widehat{G}$ (an automorphism $\alpha$ sends $\chi$ to $\chi\circ\alpha$), and any non-trivial $\chi\in\widehat{G}$ is in the orbit of one of the three weights in the above table. Moreover, for an automorphism $\alpha$ of $\mathrm{Dic}_{8}$ (viewed as an automorphism of $G$) we have that $\psi_{1}\circ\alpha\cong\psi_{1}$ and $\psi_{2}\circ\alpha\cong\psi_{2}$. Observe that for $(w_{1},w_{2})\in W_{1}\oplus W_{2}$, the condition that none of $w_{i}$ ($i=1,2)$ is zero is equivalent to the condition that $\{g\cdot(w_{1},w_{2})\mid g\in G\}$ spans $W_{1}\oplus W_{2}$ as a $K$-vector space. Therefore by Lemma 3.11 (iii), no $(w_{1},w_{2})$ with $w_{1}\neq 0$, $w_{2}\neq 0$ is contained in the common zero locus $\mathcal{V}(f\mid f\in K[W_{1}\oplus W_{2}]^{G,\chi\circ\alpha},\ \deg(f)\leq 4)$ if no such element of $W_{1}\oplus W_{2}$ is contained in $\mathcal{V}(f\mid f\in K[W_{1}\oplus W_{2}]^{G,\chi},\ \deg(f)\leq 4)$.
Note finally that for the trivial weight $(1,1,1)\in\widehat{G}$ we may take $f=1\in K[W_{1}\oplus W_{2}]^{G}$. ∎
Theorem 6.4.
Assume that $K$ contains an element of multiplicative order $8$. Then we have $\beta_{\mathrm{sep}}^{K}(\mathrm{Dic}_{8}\times\mathrm{C}_{2})=6$.
Proof.
By Theorem 2.1 we know that $\beta_{\mathrm{sep}}^{K}(\mathrm{Dic}_{8})=6$, and since $\mathrm{Dic}_{8}$ is a direct factor of $G$, we have $\beta_{\mathrm{sep}}^{K}(G)\geq\beta_{\mathrm{sep}}^{K}(\mathrm{Dic}_{8})=6$.
Now we turn to the reverse inequality. By Lemma 3.1 (iii) it is sufficient to deal with the case when $|K|$ is large enough so that we can apply Lemma 3.6. Therefore it remains to prove that $\beta_{\mathrm{sep}}(G,V)\leq 6$ where $V=W_{1}\oplus W_{2}\oplus U$. That is, we have to show that if all invariants of degree at most $6$ take the same value on $v,v^{\prime}\in V$, then $G\cdot v=G\cdot v^{\prime}$. So $v=(w_{1},w_{2},u_{\chi}\mid\chi\in\widehat{G})$ and $v^{\prime}=(w^{\prime}_{1},w^{\prime}_{2},u^{\prime}_{\chi}\mid\chi\in\widehat{G})$. It follows from Proposition 6.2 that replacing $v^{\prime}$ by an appropriate element in its $G$-orbit, we may assume that $w_{1}=w^{\prime}_{1}$ and $w_{2}=w^{\prime}_{2}$. If $w_{1}=0$ or $w_{2}=0$, then $v$ and $v^{\prime}$ belong to the same orbit by Proposition 6.1. For any $\chi\in\widehat{G}$ there exists a relative invariant $f\in K[W_{1}\oplus W_{2}]^{G,\chi^{-1}}$ with $\deg(f)\leq 4$ and $f(w_{1},w_{2})\neq 0$ by Proposition 6.3. Consider the invariant $ft_{\chi}\in K[V]^{G}$: it has degree at most $5$, and so from $(ft_{\chi})(v)=(ft_{\chi})(v^{\prime})$ we deduce $u_{\chi}=u^{\prime}_{\chi}$. This holds for all $\chi\in\widehat{G}$, thus we showed $v=v^{\prime}$, as claimed. ∎
@@ 6.1. The group $\mathrm{Dic}_{8}\times\mathrm{C}_{2}$
The other irreducible representations of $G$ are $1$-dimensional, and can be labelled by the group $\widehat{G}=\{\pm\mathrm{i},\pm 1\}\times\{\pm 1\}\leq K^{\times}\times K^{\times}$, where $\chi=(\chi_{1},\chi_{2})\in\widehat{G}$ is identified with the representation
@@
Assume that $K$ has characteristic 00. Then we have the equality $\beta(G,V)=8$ if $V$ is any of the $G$-modules
@@
Assume that $K$ has characteristic 00. For $V:=W_{1}\oplus W_{2}\oplus W_{3}$ we have the inequality $\beta_{\mathrm{sep}}(G,V)\leq 8$.
@@
$M_{2}(w)$ $:=\begin{bmatrix}x_{2}(w)y_{2}(w)&x_{1}(w)y_{1}(w)\\ x_{1}^{3}(w)y_{1}(w)&x_{2}^{3}(w)y_{2}(w)\end{bmatrix}$
$M_{3}(w)$ $:=\begin{bmatrix}x_{2}^{3}(w)y_{1}(w)&-x_{1}^{3}(w)y_{2}(w)\\ x_{1}(w)y_{2}^{3}(w)&-x_{2}(w)y_{1}^{3}(w)\end{bmatrix}.$
@@
$M_{1}(w)\cdot\begin{bmatrix}z_{1}(w)\\ z_{2}(w)\end{bmatrix}$ $=\begin{bmatrix}f_{1}(w)\\ f_{2}(w)\end{bmatrix}$
$M_{2}(w)\cdot\begin{bmatrix}z_{1}(w)\\ z_{2}(w)\end{bmatrix}$ $=\begin{bmatrix}f_{1}(w)\\ f_{3}(w)\end{bmatrix}$
$M_{3}(w)\cdot\begin{bmatrix}z_{1}(w)\\ z_{2}(w)\end{bmatrix}$ $=\begin{bmatrix}f_{4}(w)\\ f_{5}(w)\end{bmatrix}$
@@
Assume that $K$ has characteristic 00. Let $i,j\in\{1,2,3\}$, $i\neq j$, and $\chi\in\widehat{G}$. Then $\beta_{\mathrm{sep}}(G,W_{i}\oplus W_{j}\oplus U_{\chi})\leq 8$.
@@
Using the CoCalc platform [44] we verified that for a $G$-module of the form $V=W_{i}\oplus W_{j}\oplus U_{\chi}$ we have $\beta(G,V)>8$ if and only if $V$ is as in the table below:
@@
We shall show that $\beta_{\mathrm{sep}}(G,V)\leq 8$ for $V=W_{2}\oplus W_{3}\oplus U_{(\mathrm{i},-1)}$; the argument for the other $V$ in the table above is the same. Take $v=(w_{2},w_{3},u)\in V$ and $v^{\prime}=(w^{\prime}_{2},w^{\prime}_{3},u^{\prime})\in V$, and assume that $f(v)=f(v^{\prime})$ holds for all $f\in K[V]^{G}$ with $\deg(f)\leq 8$. We need to show that then $G\cdot v=G\cdot v^{\prime}$. Recall that the coordinate functions on $W_{2}$ are denoted by $y_{1},y_{2}$, on $W_{3}$ by $z_{1},z_{2}$, and on $U_{(\mathrm{i},-1)}$ by $t$. By Proposition 6.5 we know that if $w_{2}=0$, then $w^{\prime}_{2}=0$, and $(w_{3},u)$ and $(w^{\prime}_{3},u^{\prime})$ have the same $G$-orbit, Similarly, if $w_{3}=0$, then $G\cdot v=G\cdot v^{\prime}$. So we may assume that $w_{2}\neq 0$ and $w_{3}\neq 0$, and again by Proposition 6.5, by replacing $v^{\prime}$ by an appropriate element in its $G$-orbit, we may assume that $w_{2}=w^{\prime}_{2}$ and $w_{3}=w^{\prime}_{3}$. It remains to show that $u=u^{\prime}$. Denote by $tf_{1}$, $tf_{2}$, $tf_{3}$ the $G$-invariants on $V$ given in the table. All have degree less than $8$, so $(tf_{j})(v)=(tf_{j})(v^{\prime})$ holds for $j=1,2,3$. One can easily deduce from $w_{2}\neq 0$ and $w_{3}\neq 0$ that $f_{j}(w_{2},w_{3})\neq 0$ for some $j\in\{1,2,3\}$, hence $t(v)=t(v^{\prime})$, i.e. $u=u^{\prime}$. ∎
@@
Assume that $K$ has characteristic 00 and it has an element of multiplictive order $12$. Then we have the equality $\beta_{\mathrm{sep}}^{K}(\mathrm{Dic}_{12}\times\mathrm{C}_{2})=8$.
@@
Since $\mathrm{Dic}_{12}$ is a homomorphic image of $G$, we have the obvious inequality $\beta_{\mathrm{sep}}^{K}(G)\geq\beta_{\mathrm{sep}}^{K}(\mathrm{Dic}_{12})$, and by Theorem 2.1 we have $\beta_{\mathrm{sep}}^{K}(\mathrm{Dic}_{12})=8$. By Lemma 3.6 it remains to prove that $\beta_{\mathrm{sep}}(G,V)\leq 8$, where $V$ is the $G$-module $V:=W_{1}\oplus W_{2}\oplus W_{3}\oplus W_{4}\oplus U$. Let $v=(w_{1},w_{2},w_{3},w_{4},u)\in V$, $v^{\prime}=(w^{\prime}_{1},w^{\prime}_{2},w^{\prime}_{3},w^{\prime}_{4},u^{\prime})\in V$, and assume that
(16) $f(v)=f(v^{\prime})\text{ for all }f\in K[V]^{G}\text{ with }\deg(f)\leq 8.$
@@
Case 1: There exists some $i,j\in\{1,2,3\}$, $i\neq j$, such that $w_{i}\neq 0$ and $w_{j}\neq 0$.
@@
Consider the following irreducible $3$-dimensional representations of $G$:
@@
We may take $A_{(\omega,1)}:=\{s_{\omega},s^{2}_{\omega^{2}}\}$ and $A_{(\omega^{2},1)}:=\{s_{\omega^{2}},s^{2}_{\omega}\}$ (this can be verified by computing a Gröbner basis in $\mathcal{H}(G,W_{2})$). Now consider an element $f=ht_{\chi^{(1)}}\cdots t_{\chi^{(k)}}\in C$. Here $\deg(h)=2$, or $\deg(h)=4$ and $h$ is the product of two $\langle a,b,d\rangle$-invariants. After a possible renumbering we may assume that $\chi^{(i)}_{2}=-1$ for each $i=1,...,\ell$ and $\chi^{(j)}_{2}=1$ for each $j=\ell+1,...,k$. Since $h$ is a $\langle d\rangle$-invariant, $\ell=2\ell^{\prime}$ must be even. We have $f=h\prod_{i=1}^{\ell^{\prime}}(t_{\chi^{(i)}}t_{\chi^{(i+\ell^{\prime})}})\prod_{j=\ell+1}^{k}t_{\chi^{(j)}}$. So the $G$-invariant $f$ is written as a product of relative invariants of weight $\chi\in\langle(\omega,1)\rangle\cong\mathrm{C}_{3}$, where the number of factors is $k-\ell^{\prime}+1$ or $k-\ell^{\prime}+2$, depending on whether $\deg(h)=2$ or $\deg(h)=4$. On the other hand, the number of factors is at most $\mathsf{D}(\mathrm{C}_{3})=3$, since $\chi^{(1)},\dots\chi^{(k)}$ is a product-one free sequence over $\widehat{G}$. Since $\deg(f)=\deg(h)+k$, we conclude in both cases that $\deg(f)\leq 6$, and we are done. ∎
@@
$|\mathrm{Stab}_{G}(v)|=\begin{cases}3&\text{ if }x_{1}^{2}(v)=x_{2}^{2}(v)=x_{3}^{2}(v)\\ 2&\text{ if }x_{j}(v)=0\text{ for a unique }j\in\{1,2,3\}\\ 4&\text{ if }x_{j}(v)\neq 0\text{ for a unique }j\in\{1,2,3\}\\ 1&\text{ otherwise.}\end{cases}.$
@@
Case II: $(x_{1}x_{2}x_{3})(v)\neq 0$ and |$|\{x_{1}^{2}(v),x_{2}^{2}(v),x_{3}^{2}(v)\}|=2$, say $x_{1}^{2}(v)=x_{2}^{2}(v)\neq x_{3}^{2}(v)$. Similarly to Case I, by basic linear algebra we conclude from $f_{1}(v)=f_{1}(v^{\prime})$ and $f_{2}(v)=f_{2}(v^{\prime})$ that
@@
Case IV: $x_{1}^{2}(v)=x_{2}^{2}(v)=x_{3}^{2}(v)$. Applying an element of $G$ and a rescaling on $W_{1}$ (see Section 3.3) we may assume that $w_{1}=[1,1,1]^{T}$. We have the $G$-invariants

### v3

@@ 4. Some groups $G=H\times\mathrm{C}_{2}$ with $\beta_{\mathrm{sep}}^{K}(G)=\beta_{\mathrm{sep}}^{K}(H)$
These are pairwise non-isomorphic (have distinct characters). The other irreducible representations of $G$ are $1$-dimensional, and can be labelled by the group $\widehat{G}=\{\pm\mathrm{i},\pm 1\}\times\{\pm 1\}\leq K^{\times}\times K^{\times}$, where $\chi=(\chi_{1},\chi_{2})\in\widehat{G}$ is identified with the representation
@@
Assume that $K$ has characteristic 0. Then we have the equality $\beta(G,V)=8$ if $V$ is any of the $KG$-modules
@@
  • (ii)
    Note that $|\ker(\psi_{4})|=4$ and $|\ker(\psi_{j})|=2$ for $j=1,2,3$. This may be an explanation for the lack of symmetry in the roles of the four $2$-dimensional representations above.
@@
Assume that $K$ has characteristic 0. For $V:=W_{1}\oplus W_{2}\oplus W_{3}$ we have the inequality $\beta_{\mathrm{sep}}(G,V)\leq 8$.
@@
$M_{2}(w)$ $:=\begin{bmatrix}x_{2}(w)y_{2}(w)&x_{1}(w)y_{1}(w)\\ x_{1}^{3}(w)y_{1}(w)&x_{2}^{3}(w)y_{2}(w)\end{bmatrix}\qquad M_{3}(w):=\begin{bmatrix}x_{2}^{3}(w)y_{1}(w)&-x_{1}^{3}(w)y_{2}(w)\\ x_{1}(w)y_{2}^{3}(w)&-x_{2}(w)y_{1}^{3}(w)\end{bmatrix}$.
@@
$M_{1}(w)\cdot\begin{bmatrix}z_{1}(w)\\ z_{2}(w)\end{bmatrix}=\begin{bmatrix}f_{1}(w)\\ f_{2}(w)\end{bmatrix}\quad M_{2}(w)\cdot\begin{bmatrix}z_{1}(w)\\ z_{2}(w)\end{bmatrix}=\begin{bmatrix}f_{1}(w)\\ f_{3}(w)\end{bmatrix}\qquad M_{3}(w)\cdot\begin{bmatrix}z_{1}(w)\\ z_{2}(w)\end{bmatrix}=\begin{bmatrix}f_{4}(w)\\ f_{5}(w)\end{bmatrix}$
@@
Assume that $K$ has characteristic 0. Let $\in\{1,2,3\}, iji\neq j$, and $\chi\in\widehat{G}$. Then $\beta_{\mathrm{sep}}(G,W_{i}\oplus W_{j}\oplus U_{\chi})\leq 8$.
@@
Using the CoCalc platform [30] we verified that for a $KG$-module of the form $V=W_{i}\oplus W_{j}\oplus U_{\chi}$ we have $\beta(G,V)>8$ if and only if $V$ is as in the table below:
@@
We shall show that $\beta_{\mathrm{sep}}(G,V)\leq 8$ for $V=W_{2}\oplus W_{3}\oplus U_{(\mathrm{i},-1)}$; the argument for the other $V$ in the table above is the same. Take $v=(w_{2},w_{3},u)\in V$ and $v^{\prime}=(w^{\prime}_{2},w^{\prime}_{3},u^{\prime})\in V$, and assume that $f(v)=f(v^{\prime})$ holds for all $f\in K[V]^{G}$ with $\deg(f)\leq 8$. We need to show that then $G\cdot v=G\cdot v^{\prime}$. Recall that the coordinate functions on $W_{2}$ are denoted by $y_{1},y_{2}$, on $W_{3}$ by $z_{1},z_{2}$. Write $t:=t_{(\mathrm{i},-1)}$ for the coordinate function on $U_{(\mathrm{i},-1)}$. By Proposition 4.1 we know that if $w_{2}=0$, then $w^{\prime}_{2}=0$, and $(w_{3},u)$ and $(w^{\prime}_{3},u^{\prime})$ have the same $G$-orbit, Similarly, if $w_{3}=0$, then $G\cdot v=G\cdot v^{\prime}$. So we may assume that $w_{2}\neq 0$ and $w_{3}\neq 0$, and again by Proposition 4.1, by replacing $v^{\prime}$ by an appropriate element in its $G$-orbit, we may assume that $w_{2}=w^{\prime}_{2}$ and $w_{3}=w^{\prime}_{3}$. It remains to show that $u=u^{\prime}$. Denote by $tf_{1}$, $tf_{2}$, $tf_{3}$ the $G$-invariants on $V$ given in the table. All have degree less than $8$, so $(tf_{j})(v)=(tf_{j})(v^{\prime})$ holds for $j=1,2,3$. One can easily deduce from $w_{2}\neq 0$ and $w_{3}\neq 0$ that $f_{j}(w_{2},w_{3})\neq 0$ for some $j\in\{1,2,3\}$, hence $t(v)=t(v^{\prime})$, i.e. $u=u^{\prime}$. ∎
@@
Assume that $K$ has characteristic 0 and it has an element of multiplictive order $12$. Then we have the equality $\beta_{\mathrm{sep}}^{K}(\mathrm{Dic}_{12}\times\mathrm{C}_{2})=8$.
@@
Since $\mathrm{Dic}_{12}$ is a homomorphic image of $G$, we have the obvious inequality $\beta_{\mathrm{sep}}^{K}(G)\geq\beta_{\mathrm{sep}}^{K}(\mathrm{Dic}_{12})$, and by [15, Theorem 2.1] we have $\beta_{\mathrm{sep}}^{K}(\mathrm{Dic}_{12})=8$. By Lemma 2.4 it remains to prove that $\beta_{\mathrm{sep}}(G,V)\leq 8$, where $V$ is the $KG$-module $V:=W_{1}\oplus W_{2}\oplus W_{3}\oplus W_{4}\oplus U$. Let $v=(w_{1},w_{2},w_{3},w_{4},u)\in V$, $v^{\prime}=(w^{\prime}_{1},w^{\prime}_{2},w^{\prime}_{3},w^{\prime}_{4},u^{\prime})\in V$, and assume that
(6) f(v)=f(v) $f(v^{\prime})\text{ for all }f\in K[V]^{G}\text{ with }\deg(f)\leq 8$.
@@
Case 1: There exists some $\in\{1,2,3\}, iji\neq j$, such that $w_{i}\neq 0$ and $w_{j}\neq 0$.
@@
It has the following two non-isomorphic irreducible $3$-dimensional representations:
@@
[…]
@@
[…]
@@
Case II: $(x_{1}x_{2}x_{3})(v)\neq 0$ and $|\{x_{1}^{2}(v),x_{2}^{2}(v),x_{3}^{2}(v)\}|=2$, say $x_{1}^{2}(v)=x_{2}^{2}(v)\neq x_{3}^{2}(v)$. Similarly to Case I, by basic linear algebra we conclude from $f_{1}(v)=f_{1}(v^{\prime})$ and $f_{2}(v)=f_{2}(v^{\prime})$ that
@@
Case IV: $x_{1}^{2}(v)=x_{2}^{2}(v)=x_{3}^{2}(v)$. Applying an element of $G$ and a rescaling on $W_{1}$ (see (3)) we may assume that $w_{1}=[1,1,1]^{T}$. We have the $G$-invariants